# Bistable quartic soliton in saturable nonlinear media

### v2

@@ Abstract
This report presents a theoretical demonstration of a novel bistable quartic soliton (BQS) in saturable nonlinear media, specifically within a realistic dispersion-engineered ridge waveguide made of Lithium Niobate ($\text{LiNbO}_{3}$). The study employs the variational method to establish the amplitude-width relationship, indicating the coexistence of stable solitons with the same duration but differing amplitudes. Additionally, the impact of shock on the bistable soliton is examined through perturbative variational analysis, supported by numerical results. The robustness of the BQS under perturbations is further investigated via linear stability analysis.
@@
Figure 1: (GVD profile for a quasi-TE mode of the $LN-SiO_{2}$ ridge-waveguide. The red dot on the GVD profile indicates the $\lambda_{0}$ at which 3OD vanishes ($\beta_{3}=0$) refereed as the quartic point (QP) and the shaded region around QP indicates the range where the approximation $\beta_{3}\approx 0$ is valid. Inset below (left): The mode confinement at $\lambda_{0}=1.73\mu m$. Inset above (left): The front-view representation of the waveguide geometry where the parameters are $w=2.2\mu m$ and $h=2.1\mu m$. Inset above (right): The schematic representation of the full set up.
In this brief report, we investigate the existence of bistable quartic-solitons (BQSs) within a LiNbO3 based waveguide that exhibits SN. Our theoretical analysis demonstrates that these robust BQSs can arise in suitable dispersion conditions. We examine the temporal dynamics of the BQSs under shock effects, which act as perturbations specially for SDGs. Additionally, we assess the stability of these BQSs against input amplitude noise and establish a stability map across a parametric space. Our findings indicate a pathway to realizing a new class of solitary waves in mediums with SN, which could be significant for applications in fiber lasers and optical communications.
We design a realistic ridge waveguide composed of LiNbO3, featuring external bias along a transverse direction. The GVD profile for a quasi-transverse-electric (TE) mode of the extraordinary ray is illustrated in Fig. 1. To excite SN, an external electric-field bias is applied along the optic axis [21], with uniform irradiation to avoid oversaturation [20]. The waveguide cross-section, system set-up, and mode confinement are illustrated as an inset (upper and lower) in Fig. 1. The GVD-profile exhibits the quartic point (red dot) encircled by a shaded elliptical region where the approximation $\beta_{3}\approx 0$ is valid and the GVD profile can be expressed by including only $\beta_{2}$ and $\beta_{4}$, terms respectively. In the proximity of the quartic point, GVD is anomalous ($\beta_{2}<0$) and GVD curvature is negative ($\beta_{4}<0$). Exactly at the quartic point, for operating wavelength $\lambda_{0}\approx 1.73\mu$m the dispersion coefficients are calculated as $\beta_{2}=-0.014$ ps2/m, $\beta_{3}=2.51\times 10^{-5}$ ps3/m and $\beta_{4}=-1.30\times 10^{-5}$ ps4/m, respectively. Note that, for focusing nonlinearity, these are the primary conditions for the exploration of localized QS and we carefully design the waveguide structure to achieve the correct sign and curvature of the GVD at operating wavelength. The optical pulse propagation in the waveguides where the nonlinear response saturates beyond a threshold power can be modeled by the standard NLSE. Under the slowly varying envelope approximation, the complex-valued electric field satisfies the following nonlinear Schrödinger equation (NLSE) [1],
@@
Here, we perform the typical scaling $\psi\rightarrow A/\sqrt{P_{0}}$, $\tau\rightarrow(t-v_{g}/z)/t_{0}$, $\xi\rightarrow z/L_{D}$, where $L_{D}=t^{2}_{0}/|\beta_{2}|$ is the dispersion length, $\delta_{m}=\beta_{m}/(m!|\beta_{2}|t^{m-2}_{0})$, $\beta_{m}$ is the mth order dispersion coefficient at the carrier frequency $\omega_{0}$. The normalized self-steepening parameter, relevant for SDG fibers, is defined as, $\tau_{sh}=(\omega_{0}t_{0})^{-1}$ where $t_{0}$, $P_{0}$, $v_{g}$, $\gamma$ being the input pulse duration, input peak power, group velocity and nonlinear coefficient, respectively. For our proposed waveguide, the refractive index of the LN core at the operating wavelength ($\lambda_{0}\approx 1.73$ $\mu$m) is around $n_{\text{core}}=2.13$ which offers a large refractive index contrast between core and silica cladding ($n_{\text{clad}}=1.44$), thereby resulting in tight mode confinement as indicated in the inset Fig. 1. For the input pulse width $t_{0}=55$ fs, the normalized dispersion coefficients are calculated as $\delta_{4}=-0.0125$, and $\delta_{3}=0.0053$. Note, the value of $\delta_{3}$ is one order less than $\delta_{4}$. For our system, we adopt the frequently used mathematical form of the SN response [29, 30, 31, 32] as,
@@
where $\mu=\pm 1$ determines focusing($+$) or defocusing (--) nonlinearity. The saturation parameter $s$ is defined as, $s=2|\beta_{2}|(t_{0}^{2}E_{0}k_{0}r_{ij}n_{0}^{3})^{-1}$ which depends on the bias field ($E_{0}$), electro-optic coefficient ($r_{ij}$) and unperturbed index of refraction ($n_{0}$) of the crystal in use [21]. Note, being a PR crystal, LN offers additional tunability in controlling the saturable parameter $s$ through external bias field. For our proposed waveguide structure, in the proximity of quartic point, the governing equation for the field (see Eq. (1)) can be approximated as,
@@
$-qg(\tau)-\delta_{2}\partial^{2}_{\tau}g(\tau)+\delta_{4}\partial^{4}_{\tau}g(\tau)+\frac{\mu g(\tau)^{3}}{1+sg(\tau)^{2}}=0.$ (4)
@@
[…]
@@
where, $\Re$ represents the real part, $\Theta=\cosh^{-1}\zeta$ with, $\zeta=1+2i\mathcal{A}\sqrt{s}$. Optimizing the static Lagrangian by Euler-Lagrange equation $\partial L/\partial j=0$ for $j=\mathcal{A},\kappa$ we get an unique relationship between $\mathcal{A}$ and $\kappa$,
@@
Figure 2: ($a$) Bistable $\mathcal{A}$-$\kappa$ relation derived analytically (solid black line) and by solving Eq. (4) numerically (red dotted line) for $\delta_{4}=-0.0125$ and $s=0.5$. The gray line indicates the relation under the limit $s\to 0$. Plot $(b)$-$(d)$ represent the shape preserving dynamics of BQS corresponds to the points 1-3 on $\mathcal{A}$-$\kappa$ curve. Upper panels indicate the output profile and temporal phase distribution across the pulse along with spectrograme in time ($\tau$) and frequency ($\Omega/2\pi$) space. In side panel the variation of peak amplitude is demonstrated. Plot $(e)$,$(f)$ represent dynamics of QS under shock effect with the value $\tau_{sh}=0.017$ calculated for the set-up. The white dashed lines indicates the analytical prediction of the shock mediated temporal shift of QS.
@@
In Fig. 2 ($b$)-($d$), the shape-invariant propagation of QS is illustrated for three distinct points on the amplitude-width curve marked as 1- 3. Two bistable QS states are depicted with the same widths but varying peak amplitudes. The upper panel displays the shape-preserving output profile alongside a uniform temporal phase distribution, reinforcing the solitonic behavior. A side panel presents the peak amplitude variation during propagation and an XFROG-spectrogram diagram, emphasizing the QS’s robustness. Notably, the amplitude fluctuation is more significant at point 3 compared to 2, attributed to point 2’s alignment with a more precise $\mathcal{A}$-$\kappa$ relation obtained through numerical methods without any approximation.
For the sake of completeness, we further investigate the dynamics of QS under shock effect which acts as a perturbation and influences the wave. Fig. 2 ($e$)-($f$) highlight the robust propagation of QS across two branches under shock conditions, with a noted temporal position shift that characterizes the shock phenomenon [1]. A variational analysis is employed to express this temporal position shift, denoted as $\Delta\tau_{w}$, considering the shock as a perturbation.
@@
where a hyperbolic secant function is considered as an ansatz and the parameter ϝ$\digamma_{s}$ is defined as, ϝ$\digamma_{s}=\frac{1}{m}-\frac{2}{m}\frac{\ln(\sqrt{m}+\sqrt{m+1})}{m\sqrt{m}\sqrt{m+1}}-\left[\frac{\cos^{-1}(1+2m)}{2m}\right]^{2}$ with $m=s\mathcal{A}_{0}^{2}$, where $\mathcal{A}_{0}$ represents the input amplitude. In Fig. 2 ($e$)-($f$), the white dotted lines indicate the analytical prediction (based on Eq. (8)) of temporal shift which shows an excellent agreement with the full numerical result.
@@
The parameters are defined as, ~$\tilde{\nabla}=\delta_{4}\partial^{4}_{\tau}-\delta_{2}\partial^{2}_{\tau}$, $\alpha_{0}=-q+\mu g^{2}\mathcal{G}(1+\mathcal{G})$, $\alpha_{1}=\mu g^{2}\mathcal{G}^{2}$, with $\mathcal{G}=(1+sg^{2})^{-1}$. We utilize the Fourier collocation method [38] to address the eigenvalue problem and obtain the entire spectrum of the linear-stability operator $\mathcal{O}$. An eigenvalue with a positive real part, $\Re(h)>0$, signifies that the perturbed stationary solution is unstable. We compute the LSA spectra by examining how $\Re(h)>0$ varies with pulse amplitude ($\mathcal{A}$) and saturation parameter ($s$), while maintaining a constant 4OD coefficient.
Figure 3: $(a)$ Instability phase diagram in $\mathcal{A}$-$s$ space. Plot $(b)$-$(d)$ demonstrate the propagation of QS corrupted by amplitude noise along with eigen-value spectra of instability growth $\Re(h_{m})$ for $s=0.5$ and $\delta_{4}=-0.0125$. The numeric values of the pulse parameters are for point 1, $\mathcal{A}=6$, $\kappa=0.58$ $q=1.65$ (see plot $(d)$), for point 2, $\mathcal{A}=1.78$, $\kappa=0.66$, $q=0.8$ (see plot $(c)$) and for point 3, $\mathcal{A}=1$, $\kappa=0.52$ $q=0.38$ (see plot $(b)$).
@@
In summary, we demonstrate the existence of novel BQS within an optical medium characterized by SN. Through a theoretical analysis employing variational optimization, we find that unique bistable solitonic states with varying amplitude profiles yet identical widths can persist. A realistic waveguide is engineered using the LN crystal, promoting an advantageous dispersion environment for BQS excitation. A perturbative variational analysis is also employed to grasp the influence of shock on BQS. Finally, a linear stability analysis is conducted to confirm the stability of this unique structure against perturbations. Our findings enhance the understanding of self-organized temporal structures as they develop under specific dispersion environment in saturable nonlinear media, indicating potential applications in high-power laser systems and communication technologies.

### v3

@@ Abstract
This report presents a theoretical demonstration of a novel bistable quartic soliton (BQS) in saturable nonlinear media, specifically within a realistic dispersion-engineered ridge waveguide made of Lithium Niobate ($\text{LiNbO}_{3}$). The study employs the variational method to establish the amplitude-width relationship, indicating the coexistence of stable solitons with the same duration but differing amplitudes. The impact of shock on the bistable soliton is examined through perturbative variational analysis, supported by numerical results. Additionally, we examine the interaction of BQS in different regimes and analyze the formation of the bound state. The robustness of the BQS under perturbations is further investigated via linear stability analysis.
I Introduction
@@
Figure 1: (GVD profile for a quasi-TE mode of the $LN-SiO_{2}$ ridge-waveguide. The red dot on the GVD profile indicates the $\lambda_{0}$ at which 3OD vanishes ($\beta_{3}=0$) referred as the quartic point (QP) and the shaded region around QP indicates the range where the approximation $\beta_{3}\approx 0$ is valid. Inset below (left): The mode confinement at $\lambda_{0}=1.73\mu m$. Inset above (left): The front-view representation of the waveguide geometry where the parameters are $w=2.2\mu m$ and $h=2.1\mu m$. Inset above (right): The schematic representation of the full set up.
In this report, we investigate the existence of bistable quartic-solitons (BQSs) within a LiNbO3 based waveguide that exhibits SN. Our theoretical analysis demonstrates that these robust BQSs can arise in suitable dispersion conditions. We examine the temporal dynamics of the BQSs under shock effects, which act as perturbations specially for SDGs. Additionally, we investigate the interaction of BQS exhibiting oscillatory and non-oscillatory decaying tail. A force analysis ensures QS with oscillatory tail may form a bound state. Finally, we assess the stability of these BQSs against input amplitude noise and establish a stability map across a parametric space includes incident wave amplitude and strength of the saturable nonlinearity. In actual experiments such parameters can be tuned through the input intensity and external bias field. Our findings indicate a pathway to realizing a new class of solitary waves in mediums with SN, which could be significant for applications in fiber lasers and optical communications. In particular, these BQSs can be a potential candidate to build switching and logic-gate devices. Advantageous energy scaling further makes them a viable tool for bistable lasers. The tunability of the bistable curve through the saturable parameter provides additional flexibility in laser applications.
II Theory and model
[…]
The optical pulse propagation in the waveguides where the nonlinear response saturates beyond a threshold power can be modeled by the standard NLSE. Under the slowly varying envelope approximation, the complex-valued electric field satisfies the following nonlinear Schrödinger equation (NLSE) [1],
@@ II Theory and model
[…]
@@
where $\mu=\pm 1$ determines focusing(++) or defocusing (-) nonlinearity. The saturation parameter $s$ is defined as, $s=2|\beta_{2}|(t_{0}^{2}E_{0}k_{0}r_{ij}n_{0}^{3})^{-1}$ which depends on the bias field ($E_{0}$), electro-optic coefficient ($r_{ij}$) and unperturbed index of refraction ($n_{0}$) of the crystal in use [21]. Note, being a PR crystal, LN offers additional tunability in controlling the saturable parameter $s$ through external bias field. For our proposed waveguide structure, in the proximity of quartic point, the governing equation for the field (see Eq. (1)) can be approximated as,
@@
$qg(\tau)-\delta_{2}\partial^{2}_{\tau}g(\tau)+\delta_{4}\partial^{4}_{\tau}g(\tau)+\frac{\mu g(\tau)^{3}}{1+sg(\tau)^{2}}=0.$ (4)
@@
The static Lagrangian is obtained by employing the ansatz function $g(\tau)=\mathcal{A}\operatorname{sech}^{2}(\kappa\tau)$ as, $L=\int_{-\infty}^{\infty}\mathcal{L}d\tau$ which follows,
@@
where, $\Re$ represents the real part, $\Theta=\cosh^{-1}\zeta$ with, $\zeta=1+2i\mathcal{A}\sqrt{s}$. Optimizing the static Lagrangian by Euler-Lagrange equation $\partial L/\partial j=0$ for $j=\mathcal{A},\kappa$ we get an unique relationship between $\mathcal{A}$ and $\kappa$ (See details in Appendix A),
@@
Figure 2: ($a$) Bistable $\mathcal{A}$-$\kappa$ relation derived analytically (solid black line) and by solving Eq. (4) numerically (red dotted line) for $\delta_{4}=-0.0125$ and $s=0.5$. The gray line indicates the relation under the limit $s\to 0$. Plot $(b)$ represents the shape preserving dynamics of BQS corresponds to the points 1-3 on $\mathcal{A}$-$\kappa$ curve. In side panel the variation of peak amplitude is demonstrated. In plot (c) we illustrate the output profile and temporal phase distribution across the pulse along with spectrograme in time ($\tau$) and frequency ($\Omega/2\pi$) space.
@@
In Fig. 2 ($b$), the shape-invariant propagation of QS is illustrated for three distinct points on the amplitude-width curve marked as 1- 3. Two bistable QS states are depicted with the same widths but varying peak amplitudes. A side panel presents the peak amplitude variation during propagation. Notably, in plot (b), the amplitude fluctuation is more significant at point 3 compared to 2, attributed to point 2’s alignment with a more precise $\mathcal{A}$-$\kappa$ relation obtained through numerical methods without any approximation. Fig. 2 ($c$) displays the shape-preserving output profile alongside a uniform temporal phase distribution, reinforcing the solitonic behavior and an XFROG-spectrogram diagram, emphasizing the QS’s robustness. XFROG is a standard technique used to represent ultrashort pulses in time and frequency domain and is defined as the convolution $S(\tau,\omega,\xi)=|\int_{\infty}^{\infty}\psi(\xi,\tau^{\prime})\psi_{w}(\tau-\tau^{\prime})\exp(i\omega\tau^{\prime})d\tau^{\prime}|^{2}$, where $\psi_{w}$ is the reference window function.
Effect of self-steepening: The intensity-dependent index of refraction can distort an optical pulse during propagation, potentially leading to phenomena such as optical shocks due to pulse self-steepening. This effect arises where an increase in the index causes the trailing edge of the pulse to steepen, enabling the intensity to decline as rapidly as dispersion allows. This steepening resembles the development of an acoustic shock observed at the leading edge of a sound wave and occurs in materials where the peak of the pulse travels slower than its wings, allowing the trailing part to approach the peak. For the sake of completeness, we investigate the dynamics of QS under shock effect which acts as a perturbation and influences the wave. Fig. 3 (aa)-($c$) highlight the robust propagation of QS across two branches under shock conditions, with a noted temporal position shift that characterizes the shock phenomenon [1]. A variational analysis is employed to express this temporal position shift, denoted as $\Delta\tau_{w}$, considering the shock as a perturbation (See details in Appendix B).
@@
where a hyperbolic secant function is considered as an ansatz and the parameter $\digamma_{s}$ is defined as, $\digamma_{s}=\frac{1}{m}-\frac{2}{m}\frac{\ln(\sqrt{m}+\sqrt{m+1})}{\sqrt{m}\sqrt{m+1}}-\left[\frac{\cos^{-1}(1+2m)}{2m}\right]^{2}$ with $m=s\mathcal{A}_{0}^{2}$ (see details in Appendix B), where $\mathcal{A}_{0}$ represents the input amplitude. In Fig. 3 (aa)-($c$), the white dotted lines indicate the analytical prediction (based on Eq. (8)) of temporal-shift which shows an excellent agreement with the full numerical result.
Figure 3: Plot (a)(a)-$(c)$ represent dynamics of QS under shock effect with the value $\tau_{sh}=0.017$ calculated for various $s$ values for $\delta_{4}=-0.0125$. The white dashed lines indicate the analytical prediction of the shock-mediated temporal shift of QS.
Interaction of BQS: This subsection examines the BQS interaction. In contrast to PQS, the general QS displays an oscillatory tail when the propagation constant is more than the threshold value $q_{th}=\delta_{2}^{2}/4|\delta_{4}|$ [4]. Through Ritz’s optimization of the static Lagrangian, the propagation constant $q$ is expressible as a function of amplitude $\mathcal{A}$ (See details in Appendix A).
$q=\frac{3\kappa^{2}}{8}\left(\Gamma_{2}+\frac{3\Gamma_{4}\kappa^{2}}{4}\right)+\frac{\mu}{s}\left(1-\frac{3}{4\mathcal{A}^{2}s}\Re{[\Theta^{2}]}\right).$ (9)
The propagation constant $q$ of a QS may change with amplitude and can cross the threshold value $q_{th}$, leading to the emergence of an oscillatory tail. When two QSs exhibit such exponentially decaying tails, they can interact to create a bound state [38] owing to tail overlap. The relationship, including the propagation constant $q$ versus amplitude $\mathcal{A}$ and the corresponding $\mathcal{A}$-$\kappa$ dynamics, is represented in Figure 4 ($a$) for a saturation parameter $s=0.1$. Two scenarios are considered: one, for the value of propagation constant below $q_{th}$ indicated by 1, and another above it indicated by 2. The BQS generated for $q>q_{th}$ may form bound state based on the weak interaction forces induced by the tail-tail overlapping. Exploiting the concept of momentum $P=\int_{-\infty}^{\infty}\frac{i}{2}\left(\psi\psi^{*}_{\tau}-\psi^{*}\psi_{\tau}\right)d\tau$ it is possible to calculate the interaction force as, $F=\frac{dP}{d\xi}$. For exponentially decaying tail ($q<q_{th}$) the force $F$ can be calculated as (see Appendix C for details),
$F=-g_{0}^{2}n^{2}e^{i\Delta\phi}(2q-n^{2}\delta_{2}+3n^{4}\delta_{4})e^{-2n\tau_{0}}.$ (10)
For a decaying oscillatory tail, we can form $F$ as,
$\begin{split}F=g_{0}^{2}e^{i\Delta\phi}e^{-2G\tau_{0}}\left[\delta_{2}(G\cos\alpha+K\sin\alpha)^{2}-\frac{\delta_{4}}{2}[(4G^{4}-H)\right.\\ \left.-3(4G^{2}K^{2}-H)\cos 2\alpha+12GK\sqrt{H}\sin 2\alpha]-2q\cos^{2}\alpha\right]\end{split}$ (11)
where, $H=(G^{2}-K^{2})^{2}$ and $\alpha=(K\tau_{0}+\theta)$. The values of $g_{0}$ and $\theta$ can be determined by fitting the soliton tail with $g_{0}e^{-G\tau}\cos\alpha$. The force $F$ between soliton pairs can either attract or repel depending on it’s sign, with zero force suggesting possible bound state formation. As illustrated in Figure 4 ($b$), for $q<q_{th}$, in-phase solitons attract while out-of-phase solitons repel, but no bound state forms as the force decreases to zero asymptotically. Conversely, for $q>q_{th}$, as shown in Figure 4 ($c$), the force curve exhibits a fixed point where $F=0$, indicating a bound state. As predicted by the force analysis, particularly at a separation $\Delta\tau_{0}=4.6$, a bound state appears for both in-phase and out-of-phase QS pairs.
Figure 4: (a) Propagation constant $q$ as a function of $\mathcal{A}$ (solid line) and $\mathcal{A}$-$\kappa$ relation for $\delta_{2}=-1/2$, $\delta_{4}=-0.0125$ and $s=0.1$. The horizontal solid line indicates $q_{th}$. The inset displays the soliton shape at the lower and upper branches, indicated by 1 and 2, respectively. (b) Force curve as a function of separation $\Delta\tau_{0}$ for in phase ($\Delta\phi=0$) and out-of-phase ($\Delta\phi=\pi$) BQS pair in lower branch ($q<q_{th}$). In the inset we show repulsive (for $F>0$) and attractive (for $F<0$) dynamics of BQS pair for identical separation $\Delta\tau_{0}=4.6$. (c) Force curve as a function of $\Delta\tau_{0}$ for in phase ($\Delta\phi=0$) and out-of-phase ($\Delta\phi=\pi$) BQS pair in higher branch ($q>q_{th}$). The force curves exhibit a fixed point ($F=0$) at the same separation $\Delta\tau_{0}=4.6$, and leads to the formation of a bound-state.
III Stability Analysis
@@
The parameters are defined as, $\tilde{\nabla}=\delta_{4}\partial^{4}_{\tau}-\delta_{2}\partial^{2}_{\tau}$, $\alpha_{0}=-q+\mu g^{2}\mathcal{G}(1+\mathcal{G})$, $\alpha_{1}=\mu g^{2}\mathcal{G}^{2}$, with $\mathcal{G}=(1+sg^{2})^{-1}$. We utilize the Fourier collocation method [39] to address the eigenvalue problem and obtain the entire spectrum of the linear-stability operator $\mathcal{O}$. An eigenvalue with a positive real part, $\Re(h)>0$, signifies that the perturbed stationary solution is unstable. We compute the LSA spectra by examining how $\Re(h)>0$ varies with pulse amplitude ($\mathcal{A}$) and saturation parameter ($s$), while maintaining a constant 4OD coefficient.
@@
Figure 5: $(a)$ Instability phase diagram in $\mathcal{A}$-$s$ space. Plot (b)(b)-$(d)$ demonstrate the propagation of QS corrupted by amplitude noise along with eigen-value spectra of instability growth $\Re(h_{m})$ for $s=0.5$ and $\delta_{4}=-0.0125$. The numeric values of the pulse parameters are for point 1, $\mathcal{A}=6$, $\kappa=0.58$ $q=1.65$ (see plot $(d)$), for point 2, $\mathcal{A}=1.78$, $\kappa=0.66$, $q=0.8$ (see plot $(c)$) and for point 3, $\mathcal{A}=1$, $\kappa=0.52$ $q=0.38$ (see plot $(b)$).
In summary, we demonstrate the existence of novel BQS within an optical medium characterized by saturable nonlinearity. Through a theoretical analysis employing variational optimization, we find that unique bistable solitonic states with varying amplitude profiles yet identical widths can persist. A realistic waveguide is engineered using the LN crystal, promoting an advantageous dispersion environment for BQS excitation. A perturbative variational analysis is also employed to grasp the influence of the optical shock on BQS that arises due to the intensity-dependent group velocity. Further, we investigate the interaction of the BQS pair for various conditions and by exploiting the force analysis demonstrate the formation of bound state. Finally, a linear stability analysis is conducted to confirm the stability of this unique structure against perturbations. Our findings enhance the understanding of self-organized temporal structures as they develop under specific dispersion environment in saturable nonlinear media, indicating potential applications in high-power laser systems and communication technologies.